OpenStudy (samara89):

A can of spaghetti has a diameter of 12 cm and a height of 16 cm. The can contains spherical meatballs, each with a radius of 3 cm, that take up 25% of the volume. How many meatballs are in the can? Use 3.14 to approximate pi and express your answer as a whole number.

OpenStudy (samara89):

@greatlife44

OpenStudy (samara89):

I think the answer is 4 meatballs

OpenStudy (weasel123):

the can contains spherical meatballs. no duh! i thought it had triangular meatballs

OpenStudy (greatlife44):

\[\pi*r^{2}*h_{can} = \pi*(6cm)^{2}*(16cm) =576\pi~cm^3_{volume~can}\] \[576\pi~cm^{3}*0.25 = 452cm^{3}~volume~meatballs~takes~up\] \[\frac{ 4 }{ 3 }*(3cm)^3*\pi = 113~cm^{3}_[volume of meat~ball]\] \[\frac{ Volume~meat~ball~takesup }{ meat ball volume }=\frac{ 452cm^3}{ 113cm^3 } \] \[\frac{ 452 }{ 113 } = 4~meat~balls\]

OpenStudy (greatlife44):

@Samara89 please look at what I did.